OpenStudy (anonymous):

Write an expression that is equivalent to 2/3(4x + 9) 8/3x+6 8/3x+9 2/3x+6 2/3x+9

OpenStudy (lynfran):

is ur expression \[\frac{ 2 }{ 3(4x+9) }\]or\[\frac{ 2 }{ 3 }(4x+9)\]?

OpenStudy (anonymous):

23 (4x+9) @LynFran

OpenStudy (anonymous):

\[\frac{ 2 }{ 3} (4x+9)\] I mean

OpenStudy (anonymous):

@lynfran